OpenStudy (anonymous):

How do I write this in interval notation? x values between 25 and 40 x-values less than 30 or more than 35

OpenStudy (anonymous):

The question is what x has to equal in order for it to be at least 2000, and what x should be for it to be below 2100? for 130x - 2x^2

OpenStudy (anonymous):

25 ≤ x ≤ 40, x ε [ 25, 40 ] x ≤ 30 ( ∞, 30 ]

OpenStudy (anonymous):

that's not right... what about the 35?

OpenStudy (anonymous):

@YesterdayiSaidTomorow How come yesterday you were making fun of all the kids that died in the shooting?

OpenStudy (anonymous):

x(130 - 2x) >= 2000 130x - 2x^2 >= 2000 0 <= 2x^2 - 130x + 2000 0 <= x^2 - 65x + 1000 you can either graph this, or use the quadratic formula x = [65 +/- sqrt(65^2 - 4(1)(1000))] / (2 *1) x = [65 +/- sqrt(225)] / 2 x = [65 +/- 15]/2 x = 80/2 = 40 x = 50/2 = 25 x values between 25 and 40 will give areas > 2000 130x - 2x^2 < 2100 0 < 2x^2 - 130x + 2100 0 < x^2 - 65x + 1050 x = [65 +/- sqrt(65^2 - 4(1050))] / 2 x = [65 +/- sqrt(25)] / 2 x = [65 +/- 5] / 2 x = 70/2 = 35 x = 60/2 = 30 x-values less than 30 or more than 35 will give areas less than 2100

OpenStudy (anonymous):

I need to know how to put this in interval notation

OpenStudy (anonymous):

That's easy.

OpenStudy (anonymous):

Sorry I don't help people who make fun of kids who died in a school shooting.

OpenStudy (anonymous):

but it's not -infinity to 30 and 35 to infinity

OpenStudy (anonymous):

@HorribleAtMath reported for spam.

OpenStudy (anonymous):

@YesterdayiSaidTomorow reported for being an retrice

OpenStudy (anonymous):

By the way I hope you die in a school shooting. That is all.

OpenStudy (anonymous):

I didn't do anything wrong. I just asked a question and now you're spamming it with your nonsense.

OpenStudy (anonymous):

Now you're death threatening me. Reported for verbal abuse.

OpenStudy (anonymous):

Not death threatening you? More like a death wish? You didn't do anything wrong? HAHAHAHA okay because making fun of kids dying isn't wrong...

OpenStudy (anonymous):

Please get some therapy.

OpenStudy (anonymous):

Please stop and stop spamming my question. I never did any of those things.

OpenStudy (anonymous):

Stop typing. Just go away.

OpenStudy (anonymous):

Please don't lie. I guess it's your choice if you wan't to be a liar but all I'm saying as that you are cruel and completely heartless and that I everyone who saw what you were saying in the chat hopes and prays that a shooter will come to your school and gun you down.

OpenStudy (anonymous):

Can a mod come and delete this troll's messages?

OpenStudy (anonymous):

The troll is you. HEY EVERYONE @YesterdayiSaidTomorow LIKES TO LAUGH AT THE CHILDREN WHO DIED IN THE NEWTOWN SHOOTING. Do your parents hate you..just wondering?

OpenStudy (anonymous):

Look at your messages, then look at mine. Who do you think is annoying here? You. I came to ask a question and now you won't leave me alone. I feel offended and threatened by the rude material you wrote.

OpenStudy (anonymous):

You are so pathetic. Maybe if you could remember your messages on the chat box the other day...we would know who is the annoying one. If you are offended and threaded by the "rude" material I wrote, then why don't you just find a way to get me off your question...block me or something. Your decision not mine.